xXMarcelieXx:

number 2 how do i do part b for part a i got l(x,y) = 1

xXMarcelieXx:

xXMarcelieXx:

@sillybilly123

sillybilly123:

in polar, that is \(\lim\limits_{r \to 0} \cos r \), yes?

xXMarcelieXx:

yes

sillybilly123:

yep, for part A \(\nabla (\cos (x^2 + y^2) - z )_{0} = <0,0,1>\)

sillybilly123:

or \(<0,0,-1>\)

xXMarcelieXx:

okay so then

sillybilly123:

i think you have done part B. job finished.

xXMarcelieXx:

wait what im lost was that part b what u did?

sillybilly123:

what do you think Part B is and what do you think the answer to Part B is?

sillybilly123:

every derivative is a limit. if you cannot establish a limit at a point, then ...

xXMarcelieXx:

well ik its a limit

sillybilly123:

i think you have answered these questions. you have established that the limit is \(z(0) = 1\). and that the normal vector is \(<0,0,-1>\) that is job done

xXMarcelieXx:

hmm

sillybilly123:

you doubt !!

sillybilly123:

cool!!

xXMarcelieXx:

no im not doubting lol im just wondering how is that the answer

xXMarcelieXx:

werent we suppose to use some limit?

sillybilly123:

we used a limit for Part B, having switched Cartesian to Polar so we had: \(\lim\limits_{r \to 0} r = 0\)

xXMarcelieXx:

oh okay so then

sillybilly123:

pretty sure that's it all

sillybilly123:

i can set you some questions

sillybilly123:

or not

xXMarcelieXx:

ok

sillybilly123:

are you learning physics or math?

xXMarcelieXx:

math

sillybilly123:

ouch. you won't really get this stuff unless you do a physical science. unless you have a math brain, in which case you will !!

xXMarcelieXx:

lol yeh.. ive never taken physics

sillybilly123:

it's ALL physics. i'm afraid :( but you math bods make sure it works !

xXMarcelieXx:

yah lol

sillybilly123:

😁

xXMarcelieXx:

she uploaded the solution for this problem barely... idk how she got 0

xXMarcelieXx:

1 attachment
sillybilly123:

dodgy answers :) in both cases

xXMarcelieXx:

dang

sillybilly123:

Nah!!

sillybilly123:

Mmmmm. Brutal stuff :( Not sure I would agree with most of that if you need another steer, do tag @Vocaloid , the Mother Superior i will be only to happy to addd my inept input :))